If the kinetic energy of the body becomes 4th time of its initial value than the new Momentum will

Grade:11

1 Answers

R.vamshi rathod
6 years ago
as the kinetic energy increases the velocity of the body aiso increases as 1/2mvsquare=m*v= 2vsquare=v.......so th momementum of the body also increases as the mass is same...